Exercise 13-3 (Part Level Submission)

Vaughn Company began operations on January 2, 2016. It employs 10 individuals who work 8-hour days and are paid hourly. Each employee earns 9 paid vacation days and 6 paid sick days annually. Vacation days may be taken after January 15 of the year following the year in which they are earned. Sick days may be taken as soon as they are earned; unused sick days accumulate. Additional information is as follows.

Actual Hourly
Wage Rate

Vacation Days Used
by Each Employee

Sick Days Used
by Each Employee

2016

2017

2016

2017

2016

2017

$11 $12 0 8 4 5

Vaughn Company has chosen to accrue the cost of compensated absences at rates of pay in effect during the period when earned and to accrue sick pay when earned.

Collapse question part

(a)

Prepare journal entries to record transactions related to compensated absences during 2016 and 2017. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

2016

(To accrue the expense and liability for vacations)

(To accrue the expense and liability for sick pay)

(To record payment for compensated time when used by employees)

2017

(To accrue the expense and liability for vacations)

(To accrue the expense and liability for sick pay)

(To record vacation time paid)

(To record sick leave paid)
Answer #1

ANSWER:

Ref Account Tittle Debit Credit
-1 2016 Salaries and Wages Expense 7,920
Salaries and Wages Payable 7,920
To accrue the expense and liability for vacations
-2 2016 Salaries and Wages Expense 3,520
Salaries and Wages Payable 3,520
To record sick leave paid
2016 No entry,
since no vacation days were used.
To record vacation time paid
-3 2017 Salaries and Wages Expense 8,640
Salaries and Wages Payable 8,640
To accrue expense and liability for vacations
-4 2017 Salaries and Wages Expense 4,800
Cash 4,800
To record sick leave paid
2017 Salaries and Wages Expense 640
-5 Salaries and Wages Payable 7,040
-6 Cash 7,680
To record vacation time paid
employees per hour hours per day days
-1 10 * 11 * 8 * 9 = 7920 -1
-2 10 * 11 * 8 * 4 = 3520 -2
-3 10 * 12 * 8 * 9 8640 -3
-4 10 * 12 * 8 * 5 = 4800 -4
-5 10 * 11 * 8 * 8 = 7040 -5
-6 10 * 12 * 8 * 8 = 7680 -6

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************